OpenStudy (anonymous):

I need help *** if someone can get back to me you will get a medal!! really need help lost

OpenStudy (amistre64):

Post the question you need help with, and not an advertisement to "swap answers" please

OpenStudy (anonymous):

what quiz

OpenStudy (anonymous):

ok for the function 32 over x+3 find (g^-1*g(5)

OpenStudy (anonymous):

Graphing Radical Functions Quiz '

OpenStudy (amistre64):

what is the value of g(5)?

OpenStudy (anonymous):

Sorry it soupposed to be (g^-1*g)(5)

OpenStudy (anonymous):

I dont remeber that, srry I take Geometry and Id ont rmeber how to solve

OpenStudy (anonymous):

it does not say what g equals

OpenStudy (amistre64):

im pretty sure g is the name of the function they gave you

OpenStudy (amistre64):

\[g(x)=\frac{32}{x+3}\]

OpenStudy (anonymous):

ok so i multiply that funtion by itself to the -1 power?

OpenStudy (amistre64):

the -1 notation here denotes the inverse of the function .. not really a power

OpenStudy (amistre64):

do you know how to find an inverse?

OpenStudy (anonymous):

not really

OpenStudy (amistre64):

given a function \[y=\frac{32}{x+3}\] the inverse swapps x and y, so the inverse of this is \[x=\frac{32}{y+3}\]

OpenStudy (anonymous):

so now i divide?

OpenStudy (amistre64):

lets work the inverse a little more .. just so things loook normal \[x=\frac{32}{y+3}\] \[y+3=\frac{32}{x}\] \[y=\frac{32}{x}-3=g^{-1}(x)\]

OpenStudy (amistre64):

so, \[g^{-1}*g=(\frac{32}x-3)(\frac{32}{x+3})\] plug in x=5 and find the product

OpenStudy (anonymous):

4

OpenStudy (anonymous):

now what? would i times 4 by 5?

OpenStudy (amistre64):

not if your following what ive written ....

OpenStudy (anonymous):

the answer choices are 11 5 2 and 1 so ???

OpenStudy (anonymous):

what now

OpenStudy (anonymous):

sorry i am not good at this

OpenStudy (amistre64):

hmm, you sure thats a "*" in between them or is it more of an "\(^o\)"

OpenStudy (anonymous):

an o

OpenStudy (amistre64):

ahhh, that changes things, the open "o" means a composite, not a multiplication of the 2 of them

OpenStudy (amistre64):

g^(-1) o g has a property that the input is the same as the output

OpenStudy (anonymous):

could you check out the math chat box it may need to be controlled.

OpenStudy (amistre64):

it needs the adminstrators to come online and deal with the isp itself .. so that the whole school gets banned

OpenStudy (amistre64):

so if the input is 5, what is the output?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

-5

OpenStudy (amistre64):

yes, if the input is 5, the output is 5 ... its a special property of composing inverses of each other

OpenStudy (anonymous):

what school. Cant you bn them if they need

OpenStudy (anonymous):

so if there is an o don't multiply make input sam as output?

OpenStudy (amistre64):

only if the functions are inverses of each other does that work. If they are not inverses the input and output can be different

OpenStudy (anonymous):

brb domestic situation at house have to help brother

OpenStudy (amistre64):

good luck ::)